# The Disparate Benefits of Deep Ensembles

### v1

@@ Appendix E Complete Experimental Results
Table 11: PP results (accuracy and fairness violation metrics) on UTK. Models are trained on target variable age, evaluated using protected attribute race. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before PP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After PP PP-SPD ($\downarrow$) PP-EOD ($\downarrow$) PP-AOD ($\downarrow$)
@@
Table 12: PP results (accuracy and fairness violation metrics) on UTK. Models are trained on target variable gender, evaluated using protected attribute age. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before PP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After PP PP-SPD ($\downarrow$) PP-EOD ($\downarrow$) PP-AOD ($\downarrow$)
@@
Table 13: PP results (accuracy and fairness violation metrics) on UTK. Models are trained on target variable gender, evaluated using protected attribute race. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before PP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After PP PP-SPD ($\downarrow$) PP-EOD ($\downarrow$) PP-AOD ($\downarrow$)
@@
Table 14: PP results (accuracy and fairness violation metrics) on UTK. Models are trained on target variable race, evaluated using protected attribute age. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before PP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After PP PP-SPD ($\downarrow$) PP-EOD ($\downarrow$) PP-AOD ($\downarrow$)
@@
Table 15: PP results (accuracy and fairness violation metrics) on UTK. Models are trained on target variable race, evaluated using protected attribute gender. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before PP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After PP PP-SPD ($\downarrow$) PP-EOD ($\downarrow$) PP-AOD ($\downarrow$)
@@
Table 16: PP results (balanced accuracy and fairness violation metrics) on CX. Models are evaluated using protected attribute age. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before PP BAcc ($\uparrow$) SPD ($\downarrow$) BAcc ($\uparrow$) EOD ($\downarrow$) BAcc ($\uparrow$) AOD ($\downarrow$)
@@
After PP PP-SPD ($\downarrow$) PP-EOD ($\downarrow$) PP-AOD ($\downarrow$)
@@
Table 17: PP results (balanced accuracy and fairness violation metrics) on CX. Models are evaluated using protected attribute gender. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before PP BAcc ($\uparrow$) SPD ($\downarrow$) BAcc ($\uparrow$) EOD ($\downarrow$) BAcc ($\uparrow$) AOD ($\downarrow$)
@@
After PP PP-SPD ($\downarrow$) PP-EOD ($\downarrow$) PP-AOD ($\downarrow$)
@@
Table 18: PP results (balanced accuracy and fairness violation metrics) on CX. Models are evaluated using protected attribute race. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before PP BAcc ($\uparrow$) SPD ($\downarrow$) BAcc ($\uparrow$) EOD ($\downarrow$) BAcc ($\uparrow$) AOD ($\downarrow$)
@@
After PP PP-SPD ($\downarrow$) PP-EOD ($\downarrow$) PP-AOD ($\downarrow$)
@@ Appendix F Additional Investigations
This section presents additional investigations that are complementary to those presented in the main section of the manuscript. First, we investigate how the disparate benefits effect behaves for different model sizes of the individual ensemble members. Second, we conduct the same investigation on different model architectures. Third, we study whether the disparate benefits effect also occurs for heterogeneous Deep Ensembles composed of members with different model architectures. Fourth, we report an alternative approach to mitigate the negative impact on fairness due to Deep Ensembling by means of weighting individual members differently in the ensemble. Finally, we study the sensitivity of the Deep Ensemble and its individual members to the threshold used to make the prediction.
F.1 Model Size
The experiments in the main paper were conducted using ResNet50 models. In this section we investigate whether the size of the models plays a major role in determining the existence and strength of the disparate benefits effect. The results are shown in Fig. 17 - 19. As seen in the Figures, in the majority of cases the performance gains due to ensembling slightly increase for larger model classes. The fairness violations however increase to a larger degree, see e.g. Fig. 17 (a) and (b), Fig. 18 (a), (b) and (c) as well as Fig. 19 (a). Generally, we observe an increase in the magnitude of the change in fairness violations with larger model classes for all tasks that exhibit significant disparate benefits (c.f. Tab. 1).
@@ F.1 Model Size
F.2 Model Architecture
In this section we investigate the role of the specific model architecture on the existence and strength of the disparate benefits effect. The results are shown in Fig. 20 - 22. In the majority of cases, disparate benefits occur throughout all considered model architectures. Especially for EfficientNetV2-S we observe significant disparate benefits for some cases where we do not observe them in the main investigation based on ResNet50. For example for UTK, target race, group age under AOD (Fig. 21e) or CX, group race under EOD and AOD. Overall, we do not find a systematic difference of the results for different model architectures.
@@ F.3 Heterogenous Ensembles
F.4 Deep Ensemble Weighting
@@
Fairness-based weighting. Furthermore, we leveraged the information about the fairness violation of the individual members to define the weights and yield a fairer ensembling. Given a fairness violation measure $F_{n}\in[0,1]$ for each ensemble member, we define the weighting factor
@@
F.5 Thresholds Selection
Finally, we report the results of analyzing the dependency of the Deep Ensemble and individual ensemble members on selecting the threshold for prediction. When using the usual $\mathop{\mathrm{argmax}\,}$, implicitly a threshold of 0.5 is used. In the post-processing experiments, we found that applying the method even under an additional fairness constraint can improve the performance. We evaluated all trained models on their respective validation datasets. The results are depicted in Fig. 25. The results show that the Deep Ensemble is more sensitive to the threshold on the FF dataset, especially for target variable age. Regarding the CX dataset, the balanced accuracy exhibits roughly the same behavior under varying thresholds for the Deep Ensemble than for individual members. However, the spread of the optimal threshold is much smaller throughout all experiments.

### v2

@@ Appendix E Complete Experimental Results
Table 12: HPP results (accuracy and fairness violation metrics) on UTK. Models are trained on target variable age, evaluated using protected attribute race. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before HPP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After HPP HPP-SPD ($\downarrow$) HPP-EOD ($\downarrow$) HPP-AOD ($\downarrow$)
@@
Table 13: HPP results (accuracy and fairness violation metrics) on UTK. Models are trained on target variable gender, evaluated using protected attribute age. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before HPP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After HPP HPP-SPD ($\downarrow$) HPP-EOD ($\downarrow$) HPP-AOD ($\downarrow$)
@@
Table 14: HPP results (accuracy and fairness violation metrics) on UTK. Models are trained on target variable gender, evaluated using protected attribute race. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before HPP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After HPP HPP-SPD ($\downarrow$) HPP-EOD ($\downarrow$) HPP-AOD ($\downarrow$)
@@
Table 15: HPP results (accuracy and fairness violation metrics) on UTK. Models are trained on target variable race, evaluated using protected attribute age. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before HPP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After HPP HPP-SPD ($\downarrow$) HPP-EOD ($\downarrow$) HPP-AOD ($\downarrow$)
@@
Table 16: HPP results (accuracy and fairness violation metrics) on UTK. Models are trained on target variable race, evaluated using protected attribute gender. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before HPP Acc ($\uparrow$) SPD ($\downarrow$) Acc ($\uparrow$) EOD ($\downarrow$) Acc ($\uparrow$) AOD ($\downarrow$)
@@
After HPP HPP-SPD ($\downarrow$) HPP-EOD ($\downarrow$) HPP-AOD ($\downarrow$)
@@
Table 17: HPP results (balanced accuracy and fairness violation metrics) on CX. Models are evaluated using protected attribute age. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before HPP BAcc ($\uparrow$) SPD ($\downarrow$) BAcc ($\uparrow$) EOD ($\downarrow$) BAcc ($\uparrow$) AOD ($\downarrow$)
@@
After HPP HPP-SPD ($\downarrow$) HPP-EOD ($\downarrow$) HPP-AOD ($\downarrow$)
@@
Table 18: HPP results (balanced accuracy and fairness violation metrics) on CX. Models are evaluated using protected attribute gender. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before HPP BAcc ($\uparrow$) SPD ($\downarrow$) BAcc ($\uparrow$) EOD ($\downarrow$) BAcc ($\uparrow$) AOD ($\downarrow$)
@@
After HPP HPP-SPD ($\downarrow$) HPP-EOD ($\downarrow$) HPP-AOD ($\downarrow$)
@@
Table 19: HPP results (balanced accuracy and fairness violation metrics) on CX. Models are evaluated using protected attribute race. Statistics are obtained from five independent runs, and additionally over all individual ensemble members if applicable.
Before HPP BAcc ($\uparrow$) SPD ($\downarrow$) BAcc ($\uparrow$) EOD ($\downarrow$) BAcc ($\uparrow$) AOD ($\downarrow$)
@@
After HPP HPP-SPD ($\downarrow$) HPP-EOD ($\downarrow$) HPP-AOD ($\downarrow$)
@@ Appendix F Additional Investigations
This section presents additional investigations that are complementary to those presented in the main section of the manuscript. First, we analyze the complementary notion of min-max fairness. Second, we investigate how the disparate benefits effect behaves for different model sizes of the individual ensemble members. Third, we conduct the same investigation on different model architectures. Fourth, we study whether the disparate benefits effect also occurs for heterogeneous Deep Ensembles composed of members with different model architectures. Fifth, we report an alternative approach to mitigate the negative impact on fairness due to Deep Ensembling by means of weighting individual members differently in the ensemble. Finally, we study the calibration of the Deep Ensemble and its individual members and the resulting sensitivity of their threshold used to make the prediction.
F.1 Minimax Fairness
The notions of group fairness discussed throughout the paper (Eq. (2) - (4)) control for the gap between group characteristics such as their PR, TPR or FPR. Another notion often considered in recent work is minimax fairness (Martinez et al., 2020; Diana et al., 2021; Zietlow et al., 2022), where the characteristics of the worst group are of importance. For instance (Zietlow et al., 2022) showed, that the accuracy and TPR of both the minority and majority group decrease when using standard in-processing interventions in facial analysis tasks similar to FF and UTK in our experiments. Therefore, we investigate the minimax fairness impact of Deep Ensembles. Specifically, we discuss the TPR, FPR and accuracy.
The results for TPR and FPR are given in Fig. 12 - 14. We observe, that for none of the considered tasks, there as a significant negative change of the TPR due to ensembling. Similarly, we find that for none of the considered tasks, there is a significant positive change of the FPR due to ensembling, which is desired as a better classifier should have a lower FPR. The results for accuracy are given in Fig. 18 - 20. We find, that the accuracies of both groups significantly increase for all considered tasks. In sum, while we find that Deep Ensembles have a disparate benefits effect, where one group benefits more than the other, thus increases unfairness w.r.t. disparity based group fairness metrics, the predictive performances of both groups increase thus improve fairness under a minimax fairness perspective.
(a) $Y=$ age, $A=$ gender
(b) $Y=$ age, $A=$ race
(c) $Y=$ gender, $A=$ age
(d) $Y=$ gender, $A=$ race
(e) $Y=$ race, $A=$ age
(f) $Y=$ race, $A=$ gender
Figure 18: Change in Accuracy for a Deep Ensemble (10 members) on the FF dataset, Statistics are computed based on five independent runs.
(a) $Y=$ age, $A=$ gender
(b) $Y=$ age, $A=$ race
(c) $Y=$ gender, $A=$ age
(d) $Y=$ gender, $A=$ race
(e) $Y=$ race, $A=$ age
(f) $Y=$ race, $A=$ gender
Figure 19: Change in Accuracy for a Deep Ensemble (10 members) on the UTK dataset, Statistics are computed based on five independent runs.
(a) $A=$ age
(b) $A=$ gender
(c) $A=$ race
Figure 20: Change in Accuracy for a Deep Ensemble (10 members) on the CX dataset, Statistics are computed based on five independent runs.
F.2 Model Size
The experiments in the main paper were conducted using ResNet50 models. In this section we investigate whether the size of the models plays a major role in determining the existence and strength of the disparate benefits effect. The results are shown in Fig. 21 - 23. As seen in the Figures, in the majority of cases the performance gains due to ensembling slightly increase for larger model classes. The fairness violations however increase to a larger degree, see e.g. Fig. 21 (a) and (b), Fig. 22 (a), (b) and (c) as well as Fig. 23 (a). Generally, we observe an increase in the magnitude of the change in fairness violations with larger model classes for all tasks that exhibit significant disparate benefits (cf. Tab. 1).
@@ F.2 Model Size
F.3 Model Architecture
In this section we investigate the role of the specific model architecture on the existence and strength of the disparate benefits effect. The results are shown in Fig. 24 - 26. In the majority of cases, disparate benefits occur throughout all considered model architectures. Especially for EfficientNetV2-S we observe significant disparate benefits for some cases where we do not observe them in the main investigation based on ResNet50. For example for UTK, target race, group age under AOD (Fig. 25e) or CX, group race under EOD and AOD. Overall, we do not find a systematic difference of the results for different model architectures.
@@ F.4 Heterogenous Ensembles
F.5 Deep Ensemble Weighting
@@
Fairness-based weighting. Furthermore, we leveraged the information about the fairness violation of the individual members to define the weights and yield a fairer ensembling. This is similar to the approach reported in Kenfack et al. (2021), yet for neural networks as base models. Given a fairness violation measure $F_{n}\in[0,1]$ for each ensemble member, we define the weighting factor
@@
Figure 29: Expected Calibration Error (ECE) per group (group denoted by the hatches) for individual ensemble members and the Deep Ensemble for all considered target protected attribute combinations. Statistics are computed based on five independent runs.
F.6 Calibration and Threshold Selection
As elaborated in the main part of the paper, we find that the Deep Ensemble is better calibrated than individual members (Fig. 7(a)). Here we provide a more detailed analysis that looks into the decrease in ECE per protected group for each target / protected group attribute pair (task) we consider throughout our experiments. The results are provided in Fig. 29, showing that for some tasks, the ECE significantly differs per group, but the Deep Ensemble is more calibrated than individual members, regardless of the protected group attribute.
Finally, we report the results of analyzing the dependency of the Deep Ensemble and individual ensemble members on selecting the threshold for prediction. When using the usual $\mathop{\mathrm{argmax}\,}$, implicitly a threshold of 0.5 is used. In the post-processing experiments we found that applying the method even under an additional fairness constraint can improve the performance. We evaluated all trained models on their respective validation datasets. Results are depicted in Fig. 30. The results show that the Deep Ensemble is more sensitive to the threshold on the FF dataset, especially for target variable age. Regarding the CX dataset, the balanced accuracy exhibits roughly the same behavior under varying thresholds for the Deep Ensemble than for individual members. However, the spread of the optimal threshold is much smaller throughout all experiments.